Prove that the nullspace and range of a normal operator are unchanged by taking any positive power of the operator
Answer #1

sont Let A be a normal operator de Case 1: If A in non singular e » i ken (A) = likes (AK): Zero spaces - thus. we are clone- dimo I since ing (A*) I kelam) 8 dim (ing a) + dim (ker A) so we have kes(A* )=(ing A*) ken (A) == ing (A*) - Also ker LA)Cloin! Kes A L BOLT - Let OfXe Ken (A2) and y be arbitrary ☆ 0 : < Avely> : <xuls*yy ! ied An E (ing A*) ! Also for any meta- let XE Ken Cart) & y be arbituary о оо са с ескеба е 0.- Takt ly> = KAKI Ag> 5 6x е ( img x )* n(ing n) = karte) перед Di Awe know that dim ling (A))+dim ker(A)) = dim (V) this from here we can directly say that Ting (AR) = ing (A) |